OpenStudy (anonymous):

What is the graph of the inequality? x –3y ≤ 5

OpenStudy (anonymous):

first, i would put the equation into y-intercept form (y = mx+b).

OpenStudy (anonymous):

Thanks <3

OpenStudy (anonymous):

wait, what? haha the first one?

OpenStudy (anonymous):

the equation. i'd isolate the y variable in order to make it easier to graph.

OpenStudy (anonymous):

Okay :) sorry, i dont really know how to do this stuff haha

OpenStudy (anonymous):

haha no it's fine. i go to connections academy and the math classes are hard. okay, do you understand how to isolate y?

OpenStudy (anonymous):

yeah they are...expecially when they jump me a class because they think im ready....you put a 0, right?

OpenStudy (anonymous):

y=1.6 right?

OpenStudy (anonymous):

gosh i can imagine. and not exactly. subtract x from each side and then divide by -3 on each side.

OpenStudy (anonymous):

ohhh okay

OpenStudy (anonymous):

Im not sure how to divide the 3...

OpenStudy (anonymous):

Do i leave it as a fraction?

OpenStudy (anonymous):

yup. \[y \le \frac{ 1 }{ 3 }x - \frac{ 5 }{ 3 }\]

OpenStudy (anonymous):

okay, then what

OpenStudy (anonymous):

oops, but flip the inequality sign to a greater than or equal too since we divided by a negative.

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

with y-intercept form, y=mx+b, b is the number you start with on the y axis. m is your slope. so place a dot on -5/3 and go up one, to the right three and place a dot, and so on for three or four dots.

OpenStudy (anonymous):

with that information, we can eliminate two of your choices.

OpenStudy (anonymous):

so we can take out a, c?

OpenStudy (anonymous):

nope, those are the ones we leave in.

OpenStudy (anonymous):

Oh....but i thought i go over negative and up 3?

OpenStudy (anonymous):

1/3 was positive. -1x / -3 = 1/3x

OpenStudy (anonymous):

Okay i see now sorry

OpenStudy (anonymous):

it's fine! okay, do you know how to see which way to shade?

OpenStudy (anonymous):

You shade up because its positive? haha

OpenStudy (anonymous):

you do shade up haha but pick a coordinate on the graph that doesn't go through your line. (0,0) for example. plug those x and y values into your equation. if your equation is true, shade towards the point you picked, if your equation is false, shade away from the point you picked.

OpenStudy (anonymous):

OHHH hahaha i am so lost sometimes!!!

OpenStudy (anonymous):

What is the graph of the inequality? –4y ≥ 8x – 2 i do the same with this?

OpenStudy (anonymous):

no, online math is really hard because you're teaching it to yourself basically. do you understand the shading portion?

OpenStudy (anonymous):

Yes i think i do :)

OpenStudy (anonymous):

and yes. put it into y=mx+b, and shade away haha q:

OpenStudy (anonymous):

HAhaha, thanks for your time really helpful!

OpenStudy (anonymous):

anytime!